OpenStudy (anonymous):

Simplify the expressions below. Write the final product in standard form and show your work to receive full credit. 1) 2x^4(4x^2 + 3x + 1) 2) (4x – 3)(2x^2 – 7x + 1) 3) (x^2 + 4x – 3)(2x^2 + x + 6)

OpenStudy (accessdenied):

Looks like distributive property should work out... \(a(b+c) = ab + ac\) Distributive works with polynomial expressions as well (it just requires more steps to fully simplify) \(\begin{align} (a+b)(c+d) &= (a+b)c + (a+b)d \\& = ac + bc + ad + bd \end{align} \)

OpenStudy (anonymous):

Multiplication of Polynomials is what it is.

OpenStudy (anonymous):

Could you explain MY problems to me so I can hurry and understand and turn this in before I fail, please.

OpenStudy (anonymous):

Access?

OpenStudy (accessdenied):

Yeah, and distributive property works for polynomials. :) Like, for #1: \( \begin{align} \color{Green}{2x^4} (4x^2 + 3x + 1) &= \color{green}{2x^4}*4x^2 + \color{green}{2x^4} * 3x + \color{green}{2x^4} * 1 \\ &= 8x^4 * x^2 + 6x^4 * x + 2x^4 \\ &= 8x^{4+2} + 6x^{4+1} + 2x^4 \\ &= 8x^6 + 6x^5 + 2x^4 \end{align} \)

OpenStudy (anonymous):

Aright well my instructions are this --> Simplify the expressions below. Write the final product in standard form and show your work to receive full credit. SO, is that simplified above?

OpenStudy (accessdenied):

Yes. That is simplest form. The degrees are in descending order and there is nothing more that can be done to simplify.

OpenStudy (anonymous):

Alright, now I have to write that in Standard form..

OpenStudy (accessdenied):

I believe that is standard form already...

OpenStudy (anonymous):

Um.. I'm confused. It says simplify it and then write the FINAL product in standard form. So, I guess I have to write 8x^6 + 6x^5 + 2x^4 in standard form.

OpenStudy (anonymous):

but wouldn't 8x^6 + 6x^5 + 2x^4 already be in Standard form?

OpenStudy (accessdenied):

Yeah, that's what I was thinking. Standard form, to me, is just all the terms in descending degree / simplest form.

OpenStudy (anonymous):

So, what about this one? (4x – 3)(2x^2 – 7x + 1)

OpenStudy (accessdenied):

As for #2, we'd still use distributive property, just considering the binomial as one "thing" and multiplying it through. \( \begin{align} \color{green}{(4x - 3)}(2x^2 - 7x + 1) &= \color{green}{(4x - 3)}* 2x^2 + \color{green}{(4x - 3)}(-7x) + \color{green}{(4x - 3)}*1 \\ &= 2x^2 (4x) + 2x^2 (-3) + (-7x)(4x) + (-7x)(-3) + 4x - 3 \\ &= 8x^3 \underbrace{ - 6x^2 - 28x^2} \underbrace{+21x + 4x} - 3 \\ &= 8x^3 - 34x^2 + 25x - 3 \\ \end{align} \) Makes sense?

OpenStudy (anonymous):

I don't understand this at all...

OpenStudy (accessdenied):

Do you understand distributive property? That's the basis for what I'm doing in each problem. All I'm doing is taking one thing multiplied to the other and distributing it through to each term in the other, and then it's just simplifying from there.

OpenStudy (anonymous):

Not really.. :/

OpenStudy (anonymous):

Sorry if I'm bothering you, I just want to understand it, not just get the answer, you know?

OpenStudy (accessdenied):

You're not bothering me. Of course, that's completely understandable. I'm just thinking of how to explain it. :) The basic idea of distributive property is: \(a(b+c) = ab + ac\). Of course, this will apply if \(a\) is anything being multiplied to any number of terms. \(a(b+c+d) = ab + ac + ad\) Say, a=5, b=3, and c=2 5(3+2) = 5(5) = 25 by simplification 5(3+2) = 5(3) + 5(2) = 15 + 10 = 25 by distributive So it does seem that it works out.

OpenStudy (accessdenied):

So, looking back at what I was doing earlier, that green term is just the "a" in the distributive property.

OpenStudy (anonymous):

The variables confuse me, I think. Because I understand that one completely.

OpenStudy (anonymous):

You left didn't you?....

OpenStudy (accessdenied):

The \(a\), \(b\), and \(c\) just represent anything. That's basically the idea when writing the property with those variables. It just gives us a model No. It might say I left for some reason, but I won't leave unless you tell me to or I actually "have" to. :P

OpenStudy (anonymous):

(x^2 + 4x – 3)(2x^2 + x + 6) I would say that this would be 4x^4 - 3 * 3x^2 + 6.

OpenStudy (accessdenied):

hmm... First step is \( \begin{align} \color{green}{(x^2 + 4x - 3)}(2x^2 + x + 6) = \color{green}{(x^2 + 4x - 3)}(2x^2) + \color{green}{(x^2 + 4x - 3)}(x) + \color{green}{(x^2 + 4x - 3)}(6) \end{align} \) Then we have to distribute one more time.

OpenStudy (accessdenied):

Oh, that kind of went off the page \[\begin{align} &\color{green}{(x^2 + 4x - 3)}(2x^2 + x + 6) \\ &= \color{green}{(x^2 + 4x - 3)}(2x^2) + \color{green}{(x^2 + 4x - 3)}(x) + \color{green}{(x^2 + 4x - 3)}(6) \end{align} \]

OpenStudy (anonymous):

So, that's it?

OpenStudy (accessdenied):

We'll have to use distributive one more time with the 2x^2, x, and 6. Then we simplify exponents and add the like terms.

OpenStudy (anonymous):

I've been stuck on these 3 questions for 3 hours...

OpenStudy (anonymous):

I still don't understand WHEN to distribute and where.

OpenStudy (accessdenied):

It's a property that applies whenever we multiply two things where one has addition/subtraction. If we have this case anywhere, we can always apply distributive property. We have: (x^2 + 4x - 3)(2x^2) The first factor is a polynomial with three terms, which are all connected by addition/subtraction. The second is a single term. This is just like the "a" in the original thing. So, we can just multiply 2x^2 to each part of the first polynomial.

OpenStudy (accessdenied):

x^2 (2x^2) + 4x (2x^2) - 3 (2x^2) = 2x^4 + 8x^3 - 6x^2

OpenStudy (anonymous):

So, you're multiplying each part with the last part...?

OpenStudy (accessdenied):

Yes. Similarly, (x^2 + 4x - 3)(x) Take the x to each term in the first factor x^2 (x) + 4x (x) - 3 (x) = x^3 + 4x^2 - 3x

OpenStudy (anonymous):

Now explain it like that without doing it with the equation thing so I can read it better.

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

(x2 + 4x – 3)(2x2 + x + 6)

OpenStudy (anonymous):

(x^2 + 4x – 3)(2x^2 + x + 6)

OpenStudy (accessdenied):

okay (i just use the equation thing to color the terms I'm distributing, I guess it didn't help then. :P) (x^2 + 4x - 3)(6) multiply 6 to each term inside the polynomial = x^2 (6) + 4x (6) - 3 (6) = 6x^2 + 24x - 18

OpenStudy (anonymous):

Why wouldn't this one be 4x^2 - 3 * 3x^2 + 6.

OpenStudy (anonymous):

Write the whole problem like this..(x^2 + 4x – 3)(2x^2 + x + 6)

OpenStudy (anonymous):

Then go from there showing me each step I have to do..

OpenStudy (accessdenied):

(x^2 + 4x - 3)(2x^2 + x + 6) Multiply the (x^2 + 4x - 3) to each term. = (x^2 + 4x - 3)(2x^2) + (x^2 + 4x - 3)(x) + (x^2 + 4x - 3)(6) Distribute in each multiplication individually: * 2x^2 into (x^2 + 4x - 3) * x into (x^2 + 4x - 3) * 6 into (x^2 + 4x - 3) = (x^2 (2x^2) + 4x(2x^2) + (-3)(2x^2)) + (x^2(x) + 4x(x) + (-3)(x)) + (x^2(6) + 4x(6) + (-3)(6)) Simplify the products by multiplying = 2x^4 + 8x^3 - 6x^2 + x^3 + 4x^2 - 3x + 6x^2 + 24x - 18 Combine like-terms (I'll reorder it to get the like-terms together first.) = 2x^4 + (8x^3 + x^3) + (-6x^2 + 4x^2 + 6x^2) + (-3x + 24x) - 18 = 2x^4 + 9x^3 + 4x^2 + 21x - 18

OpenStudy (accessdenied):

Many of these do sometimes take quite a bit of work to actually push through. This last one is a good example.

OpenStudy (anonymous):

I see.

OpenStudy (anonymous):

So, this simplified would be all of that?

OpenStudy (accessdenied):

The simplified form would be: 2x^4 + 9x^3 + 4x^2 + 21x - 18 Also, just curious, what class are you doing this in? Or at least, what is the topic of the section you're doing this stuff in?

OpenStudy (anonymous):

Multiplication of Polynomials - Algebra 1.

OpenStudy (accessdenied):

Ah, okay. Makes sense then. Thanks. :) Do you kind of understand what I'm doing to simplify in these now?

OpenStudy (anonymous):

I do, but I don't. Could you do the last one again without the equation thing and without detailing it. Just show me how you would simplify it and get it in standard form please.

OpenStudy (anonymous):

My assignment is this. --> http://gyazo.com/88fb14b23677d4f112bb959accd5a8b5

OpenStudy (accessdenied):

(x^2 + 4x - 3)(2x^2 + x + 6) = (x^2 + 4x - 3)(2x^2) + (x^2 + 4x - 3)(x) + (x^2 + 4x - 3)(6) = (x^2 (2x^2) + 4x(2x^2) + (-3)(2x^2)) + (x^2(x) + 4x(x) + (-3)(x)) + (x^2(6) + 4x(6) + (-3)(6)) = 2x^4 + 8x^3 - 6x^2 + x^3 + 4x^2 - 3x + 6x^2 + 24x - 18 = 2x^4 + (8x^3 + x^3) + (-6x^2 + 4x^2 + 6x^2) + (-3x + 24x) - 18 = 2x^4 + 9x^3 + 4x^2 + 21x - 18 idk, I think the raw work itself is harder to follow personally.

OpenStudy (anonymous):

Did you click the link?

OpenStudy (accessdenied):

Yeah, I see it. Would you want me to [try to] explain how to do the last two? :) The last two are much more straight-forward than this one.

OpenStudy (anonymous):

Sure. :)

OpenStudy (anonymous):

You're gone.

OpenStudy (accessdenied):

Ugh, chrome crashes again. :\ We are writing an expression for area of a rectangle, which is A=bh. We are given that b=2x-4 and h=x+5 (Given in the figure) Thus, by substituting the new information, we get: A = (2x-4)(x+5) We are asked to then write it in simplified form. We may distribute twice and add the like terms to simplify. I will choose to distribute (2x-4) A = (2x-4)(x) + (2x-4)(5) = 2x(x) - 4(x) + 2x(5) - 4(5) = 2x^2 - 4x + 10x - 20 = 2x^2 + 6x - 20

OpenStudy (accessdenied):

I'm still here, not sure why it makes me appear gone after chrome crashes.

OpenStudy (anonymous):

I see. There you are! :D

OpenStudy (accessdenied):

Yeah, hello. :P You got the last post with the explanation?

OpenStudy (anonymous):

Yeah. Write a simplified polynomial expression to represent the area of the rectangle below. So, (2x-4)(x) + (2x-4)(5) = 2x(x) - 4(x) + 2x(5) - 4(5) = 2x^2 - 4x + 10x - 20 = 2x^2 + 6x - 20 is a simplified expression?

OpenStudy (accessdenied):

Yeah. It is in descending degrees / the powers of each x are from greatest-to-least., and I cannot simplify anything in it further.

OpenStudy (anonymous):

So, (2x-4)(x) + (2x-4)(5) = 2x(x) - 4(x) + 2x(5) - 4(5) = 2x^2 - 4x + 10x - 20 = 2x^2 + 6x - 20 would be my answer? for num. 4?

OpenStudy (anonymous):

Standard form means from greatest to least exponents right?

OpenStudy (accessdenied):

Yeah. (I'd also include the A=(2x-4)(x+5) as the first line, though)

OpenStudy (anonymous):

ALright. Now Num. 5..

OpenStudy (anonymous):

x-3 .. that's all it gives me..

OpenStudy (accessdenied):

Okay, #5 We have a square tile. The actual important shape is just this:  Area of a square is A = bh, or A=s^2 (since all sides are equal, b and h are the same). So, since s=(x+3), we can then write it as A = (x+3)^2 Of course, the exponent just means to multiply it to itself. A = (x+3)(x+3) Hmm.. so, you want to try this one? To try out distributive property here. :)